OpenStudy (31356):

A ball has a diameter of 14 inches. Which choice is closest to its surface area? Use 3.14 to approximate pi. MEDAL REWARDED!!

OpenStudy (31356):

A) 154 B)615 C)2462 D)1436

OpenStudy (31356):

@thomaster @jigglypuff314 @jdoe0001 @turnhue

OpenStudy (31356):

@Destinymasha

OpenStudy (31356):

@tomhue

OpenStudy (31356):

Help ASAP

OpenStudy (31356):

@whpalmer4

OpenStudy (31356):

@jim_thompson5910

OpenStudy (anonymous):

Surface area of a sphere has the formula A = 4*pi*r^2 A = 4*3.14*(14/2)^2 A = 4*3.14*7^2 Now you can simplify for the surface area.

OpenStudy (31356):

Thanks!